Prob. 7: Calculate V. in the circuit shown below. 492 + V. - 3 62
Prob. 8: Find R for the shown circuit. eq 320 320
Answer #1


Prob ? :- Consider the circuit 42 tv- 16v br we can do this problem by two method Method - (By voltage division Rule) AccordiRule we so by voltage division can write Vo = 16x4 =6.4v 4+6 Method -2 (By kvL) + Vo 42 16v & bor -16+41 +61 - 0 101 = 16 I=1prob. 8. Consider the circuit Req, borbe Z 2r 322 Step-1 Add bar & 2r at the extreme right side since they are in Series so tStep-2 Now, &r 420 are in parallel. so, on adding Parallely giver. +- + ( Let the combined resistance be R) R = 2x8 = 1.6 2+Step-3 Now, we can add be & 1.6 directly since they are in series to find Req=(6+1.6) : Req=17.6 Results Vo = 6.40 Reg = 7.6.I lope it find useful to you. you can Comment in case of any doubter related to this. Thank you. .